The perimeter of a rectangle is 18 feet, and the area of the rectangle is 20 square feet. What is the length of the rectangle?
Answer by checkley77(12569)

18=2L+2W
20=LW OR L=20/W
18=2*20/W+2W
18=40/W+2W
18=(40+2W*W)/W
18=(40+2W^2)/W
40+2W^2=W*18
40+2W^2=18W
2W^2-18W+40=0
2(W^2-9W+20)=0
2(W-5)(W-4)=0
W-5=0
W=5 ANS. L=20/5=4 ANS.
W-4=0
W=4 ANS. L=20/4=5 ANS.
PROOFS:
20=4*5
20=20
18=2*4+2*5
18=8+10
18=18